OpenStudy (gabylovesyou):

What is the value of the y variable in the solution to the following system of equations? 5x + 3y = 7 3x - 5y = -23

OpenStudy (gabylovesyou):

@satellite73 @rajathsbhat

OpenStudy (anonymous):

multiply the first equation by 3 and the second one by 5 and subtract one from the other.

OpenStudy (gabylovesyou):

is it 4?

OpenStudy (gabylovesyou):

@hba

OpenStudy (hba):

Well there are two methods we can deal these sort of linear equations with two variables with 1) Substitution 2)Elimination Which one would you prefer ?

OpenStudy (gabylovesyou):

1

OpenStudy (hba):

Well in the process of substitution we pick one equation and then substitute it into the other 5x + 3y = 7-----(1) 3x - 5y = -23----(2) Lets pick up eq(1) =) 5x + 3y = 7 1)subtract 3y by both sides.

OpenStudy (gabylovesyou):

5x = 7-3y

OpenStudy (hba):

Now divide the whole equation by 5

OpenStudy (gabylovesyou):

do you need the solution for variable y ?

OpenStudy (hba):

Just do as i say :) I am going towards that :)

OpenStudy (hba):

And Yes we need the solution to the y variable :)

OpenStudy (gabylovesyou):

idk ;/

OpenStudy (hba):

5x = 7-3y Divide both sides by 5

OpenStudy (hba):

@Gabylovesyou I know you have a busy life Please respond......

OpenStudy (gabylovesyou):

5x / 5 = 7 - 3y / 5 x = idk

OpenStudy (hba):

Yeah right :) \[x=\frac{ 7-3y }{ 5}---(3)\] Let it be eq (3) Now put Eq(3) or x in Eq(2) 3x - 5y = -23 \[3(\frac{ 7-3y }{ 5 })-5y=-23\] Solve for y :)

OpenStudy (gabylovesyou):

ok hold on

OpenStudy (gabylovesyou):

y = 110/17

OpenStudy (hba):

Do it again

OpenStudy (gabylovesyou):

-.- idk

OpenStudy (gabylovesyou):

4?

OpenStudy (hba):

Do it again _-_

OpenStudy (gabylovesyou):

1?

OpenStudy (hba):

Yeah correct y=4

OpenStudy (gabylovesyou):

ohh lol yaaay

OpenStudy (hba):

Good going (y)